### v1

# Thermal dilepton polarization and dynamics of the QCD plasma in relativistic heavy-ion collisions

Abstract
We present the first theoretical study of the polarization of lepton pairs produced in $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV Pb+Pb collisions at the LHC, using next-to-leading order (NLO) dilepton emission rates. These calculations employ a multi-stage framework to simulate the evolution of relativistic heavy-ion collisions, and to explore the sensitivity of polarization to early times. It is found that the intermediate invariant-mass dileptons are indeed probes of the thermal equilibration process, and go beyond the reach of hadronic observables. We compute the polarization anisotropy coefficient obtained with LO dilepton rates, and show that the LO and NLO results differ radically, both in trend and in magnitude, at low and intermediate lepton pair invariant masses.
Introduction. Relativistic heavy-ion collisions produce the quark-gluon plasma (QGP), an exotic state of matter described by quantum chromodynamics Jacak and Muller (2012). As the QGP expands and cools over the duration of a collision between two nuclei, the quarks can emit electromagnetic (EM) radiation which escapes the strongly interacting medium Gale et al. (2022); Paquet et al. (2016); van Hees et al. (2011); Gale and Haglin (2003); Gale and Kapusta (1987); van Hees and Rapp (2008). Therefore, photon and dilepton measurements have the potential to provide direct information about the hot medium, such as electric conductivity Rapp (2024), chemical equilibrium Wu et al. (2024); Garcia-Montero et al. (2024); Coquet et al. (2021), magnetic fields Wang et al. (2020); Wang and Shovkovy (2021, 2022); Kimura et al. (2024), and most notably, early-stage temperature Arnaldi et al. (2009a); Churchill et al. (2024a, b). Experimentally, EM probes have been recently measured in the Large Hadron Collider (LHC) Acharya et al. (2023) and the Relativistic Heavy-Ion Collider (RHIC) Abdulhamid et al. (2023); Adamczyk et al. (2015); STA (2024).
The $p_{{}_{\rm T}}$-spectrum of direct photons is sensitive to the local hydrodynamical flow, and can therefore inform the theoretical modeling Gale et al. (2022). In contrast, the invariant mass spectrum of dileptons — a complementary probe — is impervious to flow effects. For invariant mass $M\gtrsim 1$ GeV, an important source of dileptons is quark-antiquark annihilation into a virtual photon (labeled from hereon LO, a leading-order contribution) Arnaldi et al. (2009b); STA (2024); Rapp (2013)   [^1] This channel is also included in the Drell-Yan process where initial quarks are sourced from the nPDFs of the colliding ions. , yielding thermal, and also pre-equilibrium dileptons Coquet et al. (2021); Wu et al. (2024); Garcia-Montero et al. (2024). However, corrections from the strong interaction occur in the intermediate mass range (IMR), $1~{}\rm GeV\lesssim M\lesssim 3$ GeV, and even become dominant in the low mass range (LMR): $M\lesssim 1$ GeV. For completeness, one also defines a high mass range (HMR) as $M\gtrsim 3$ GeV. We will focus on the thermal mechanism in what follows.
Because of the gluon’s abundance in a hot QGP medium, new channels open up at next-to-leading order (NLO) which includes: Compton scattering ($g\,q\to\gamma^{*}\,q$ and $g\,\bar{q}\to\gamma^{*}\,\bar{q}$), as well as modified annihilation ($q\,\bar{q}\to\gamma^{*}\,g$ and $q\,\bar{q}\,g\to\gamma^{*}$, with the latter being kinematically suppressed for $M\simeq 0$). Note that the above “real” gluon emissions need to be combined with the “virtual” (1-loop) corrections to $q\,\bar{q}\to\gamma^{*}$ to obtain a finite result at strict NLO Laine (2013); Jackson (2019). As $M\to 0$, the strict expansion becomes invalid due to the physical necessity of thermal screening and the Landau-Pomeranchuck-Migdal (LPM) effect. These are addressed in an effective theory (for small $M$) which resums naively higher-order terms Arnold et al. (2001a, b); Aurenche et al. (2002a, b). Both regimes can be systematically combined in the spectral function obtained in the thermal field theory, to cover both small and large $M$ Ghisoiu and Laine (2014); Jackson and Laine (2019); Ghiglieri and Laine (2022). Together, the NLO and LPM contributions  [^2] From this point on, we consider the “NLO contributions” to include the LPM resummation for small $M$ as in Refs. Ghisoiu and Laine (2014); Jackson and Laine (2019). are found to qualitatively enhance the observed thermal dilepton spectrum in the LMR, as well as in the IMR Churchill et al. (2024a, b, c).
In addition to mass and $p_{{}_{\rm T}}$ spectra, the polarization of EM radiation provides another unique tool to study medium properties. For example, the angular distribution of the thermal lepton pair in the rest frame of the virtual photon Speranza et al. (2018); Wei and Yan (2024); Seck et al. (2023)  [^3] Also the centre-of-mass frame of the dilepton. is expected to be sensitive to plasma anisotropy Coquet et al. (2024); the same can be said for the polarization of real photons Hauksson and Gale (2024).
This work aims to provide a realistic study of the polarization of thermal dileptons produced in heavy-ion collisions at the LHC, using state-of-the-art multi-stage modeling calibrated to reproduce hadronic observables at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV Wu et al. (2024) together with strong coupling corrected EM production rates. We introduce the polarization coefficient, and highlight its explicit dependence on the local flow and the thermal spectral densities. We show how dilepton polarization in the IMR has the potential to reveal the importance of NLO contributions, and thus the role of gluons in the medium. We also show results for the LMR, with the understanding that processes involving composite hadrons will play a crucial role there Rapp (2013), while being subdominant in the IMR. We also estimate polarization results for the elusive pre-equilibrium phase of relativistic heavy-ion collisions.
Theoretical Setup. We consider a fluid cell in thermodynamic equilibrium, characterized by a four-velocity $u^{\mu}$ and a temperature $T$  [^4] The temperature is defined in the fluid rest frame. . The differential emission rate $R_{\ell\bar{\ell}}$ of dileptons is Kapusta and Gale (2023)
$E_{+}E_{-}\frac{{\rm d}R_{\ell\bar{\ell}}}{{\rm d}^{3}\bm{p}_{+}{\rm d}^{3}\bm{p}_{-}}\;=\;-\frac{2e^{4}}{(2\pi)^{6}}\frac{\sum_{i}^{N_{\rm f}}Q_{i}^{2}}{K^{4}}\,L^{\mu\nu}\,\rho_{\mu\nu}\,f_{{}_{\rm B}}(\omega)\;.$ (1)
Here, $K=P_{+}+P_{-}$ is the four-momentum of the virtual photon, $f_{{}_{\rm B}}(\omega)\equiv\frac{1}{e^{\omega/T}-1}$ is the Bose-Einstein distribution, and $eQ_{i}$ is the charge of quark flavor $i\,$. In this work $N_{\rm f}=3$. The leptonic tensor is $L^{\mu\nu}\equiv P_{+}^{\mu}P_{-}^{\nu}+P_{-}^{\mu}P_{+}^{\nu}-g^{\mu\nu}(P_{+}\cdot P_{-}+m_{\ell}^{2})\,$, and the photon spectral function $\rho_{\mu\nu}$ is a function of the two invariants $\omega\equiv u\cdot K$ and $k\equiv\sqrt{\omega^{2}-K^{2}}\,$ (which correspond to the energy and momentum of the virtual photon in the medium’s rest frame), formally obtained by analytic continuation of the Euclidean correlator $G_{\mu\nu}$ :
$\rho_{\mu\nu}(\omega,\bm{k})$$\equiv$ $-\,{\rm Im}\bigg{[}\int_{0}^{1/T}{\rm d}\tau\,e^{i\omega_{n}\tau}G_{\mu\nu}(\tau,\bm{k})\bigg{]}_{i\omega_{n}\to\omega+i0^{+}}\,,$
$G_{\mu\nu}(\tau,\bm{k})\;$ $=$ $\;\int{\rm d}^{3}\bm{x}\,e^{-i\bm{k}\cdot\bm{x}}\big{\langle}J_{\mu}(\tau,\bm{x})J_{\nu}(0,\bm{0})\big{\rangle}_{T}\,,$ (2)
where $J_{\mu}=\bar{\psi}\gamma_{\mu}\psi$ is the EM current, $\langle...\rangle_{T}$ denotes the thermal average, and $\tau$ is the imaginary-time Laine and Vuorinen (2016).
As the medium velocity $u^{\mu}$ specifies a preferred direction, we can decompose $\rho^{\mu\nu}$ into longitudinal and transverse parts, respectively as Weldon (1982)
$\rho_{{}_{\rm L}}\;\equiv\;-\frac{K^{2}\;\rho_{\mu\nu}u^{\mu}u^{\nu}}{(u\cdot K)^{2}-K^{2}}\,,\quad\rho_{{}_{\rm T}}\;\equiv\;\frac{\rho^{\mu}_{\mu}-\rho_{{}_{\rm L}}}{2}\;.$ (3)
It is also useful to define the “vector channel” and the polarization difference, namely
$\rho_{{}_{\rm V}}\;\equiv\;\rho^{\mu}_{\mu}=\rho_{{}_{\rm L}}+2\rho_{{}_{\rm T}}\,,\quad\rho_{{}_{\Delta}}\;\equiv\;\rho_{{}_{\rm T}}-\rho_{{}_{\rm L}}=\frac{\rho_{{}_{\rm V}}-3\rho_{{}_{\rm L}}}{2}.$ (4)
From Eq. (1), one can integrate the relative momentum of the $\ell\bar{\ell}$-pair to obtain the dilepton production rate in terms of the 4-momentum of the virtual photon alone,
$\frac{{\rm d}R_{\ell\bar{\ell}}}{{\rm d}^{4}K}\;=\;\frac{2\alpha_{\rm em}^{2}}{9\pi^{2}}\ \frac{\sum_{i}^{N_{\rm f}}Q_{i}^{2}}{K^{2}}\,B\bigg{(}\frac{m_{\ell}^{2}}{K^{2}}\bigg{)}\,\rho_{{}_{\rm V}}\,f_{{}_{\rm B}}(\omega)\;,$ (5)
where $B(\xi)\equiv(1+2\xi)\sqrt{1-4\xi}$ is a phase-space factor. As the equilibrium rate in Eq. (5) is proportional to $\rho_{{}_{\rm V}}$, observables such as dilepton invariant mass spectra and elliptic flow cannot differentiate between $\rho_{{}_{\rm T}}$ and $\rho_{{}_{\rm L}}$.
Non-perturbative constraints on the spectral function can be obtained from lattice QCD simulations, where $G_{\mu\nu}(\tau,\bm{k})$ is measured at fixed momenta $k=2\pi\,n/(aN_{s})$ where $n$ is an integer, $a$ is the lattice spacing and $N_{s}$ is the number of spatial sites. In practice, the analytic continuation to real frequencies is fraught with difficulty Meyer (2011). Yet significant progress has been made, first focusing on $\rho_{{}_{\rm V}}$ Ghiglieri et al. (2016). However, the vector channel is insensitive to thermal physics owing to the large ultraviolet vacuum component of the photon self-energy. For this reason, it was suggested to consider the correlator $\rho_{\Delta}$ which vanishes in vacuum and is highly suppressed for large $\omega$ Brandt et al. (2018); Cè et al. (2022); Meyer et al. (2023). Recently, Ref. Ali et al. (2024) of the HotQCD collaboration obtained estimates for $\rho_{{}_{\Delta}}(\omega,k)$ for $2$+$1$ flavours  [^5] Although these results were at a pion mass $m_{\pi}\simeq 320$ MeV and not continuum extrapolated, the lattice was quite large with $N_{s}=96$ and $a\simeq 7$ GeV-1. at $T\simeq 220$ MeV with spatial momenta ranging from $k\simeq 0.5$ to $1.4$ GeV. Ubiquitously, $\rho_{{}_{\Delta}}$ decreases as a function of $\omega\,$, from positive at $\omega=k$ and then quickly becoming negative for $\omega>k\,$. This is consistent with perturbation theory, where both $\rho_{{}_{\rm T}}$ and $\rho_{{}_{\rm L}}$ have been worked out at strict NLO Laine (2013); Jackson (2019), and also generalized to finite baryon density Jackson (2022); Churchill et al. (2024b).
Complementary to the lattice agenda, $\rho_{{}_{\Delta}}$ is also relevant for heavy-ion experiments, when it comes to the angular distribution of the final $\ell\bar{\ell}$-pair. The latter observable, in the rest frame of the virtual photon, discriminates between the longitudinal and transverse polarization. This angular distribution can be parametrized as
$\frac{{\rm d}N}{{\rm d}^{4}K\,{\rm d}\Omega_{\ell}}$ $\propto$ $1+\lambda_{\theta}\cos^{2}\theta_{\ell}+\lambda_{\phi}\sin^{2}\theta_{\ell}\cos 2\phi_{\ell}$ (6)
$+$ $\lambda_{\theta\phi}\sin 2\theta_{\ell}\cos\phi_{\ell}+\lambda_{\phi}^{\perp}\sin^{2}\theta_{\ell}\sin 2\phi_{\ell}$
$+$ $\lambda^{\perp}_{\theta\phi}\sin 2\theta_{\ell}\sin\phi_{\ell}\;.$
A common choice of the $\gamma^{*}$ rest frame is the so-called helicity (HX) frame where the $z$-axis aligns with the momentum of the virtual photon Faccioli et al. (2010). Here, polar angles $\Omega_{\ell}=(\theta_{\ell},\phi_{\ell})$ are defined in a way such that $\ell^{+}$ in the HX frame has the 3-momentum $\bm{l}_{+}=l\,(\sin\theta_{\ell}\cos\phi_{\ell},\sin\theta_{\ell}\sin\phi_{\ell},\cos\theta_{\ell})$. By performing a change of variables and integrating out the radial part of dilepton relative momentum in Eq. (1), we arrive at an expression for the polarization coefficient $\lambda_{\theta}$ of a single fluid cell at temperature $T$ :
$\lambda_{\theta}\;=\;\frac{3(\chi-\frac{1}{3})\left(1-4\xi\right)\rho_{{}_{\Delta}}}{\frac{4}{3}(1+2\xi)\rho_{{}_{\rm V}}-(\chi-\frac{1}{3})(1-4\xi)\rho_{{}_{\Delta}}}\;,$ (7)
where $\xi\equiv\frac{m_{\ell}^{2}}{K^{2}}$ and $\chi\equiv\frac{(\bm{u}_{*}\cdot\hat{k})^{2}}{\bm{u}^{2}_{*}}\,$ with $\bm{u}_{*}$ denoting the fluid velocity viewed in the HX frame. It is related to $\bm{u}$ in the lab frame by a Lorentz boost
$\bm{u}_{*}=\bm{u}+\left(\frac{\omega}{M}-1\right)(\bm{u}\cdot\hat{k})\hat{k}-\frac{u_{0}\bm{k}}{M}.$ (8)
One can work out the other coefficients from Eq. (6), in a similar manner. Nevertheless, $\lambda_{\theta}$ is the only coefficient that does not vanish for a medium at rest, making it an indicator of $\rho_{{}_{\Delta}}$ that is less sensitive to local flow conditions. In this letter, we focus on $\lambda_{\theta}$.
Figure 1: Approximation to $\lambda_{\theta}$ from the ratio of spectral functions in Eq. (9), as a function of the invariant mass $M$ with $k$ fixed by Eq. (10). The resummed NLO result is shown for fixed coupling, $\alpha_{s}=\{0.1,0.2,0.3\}$.
We note that Eq. (7) is closely connected with $\rho_{{}_{\Delta}}$ Seck et al. (2023). Indeed, for a plasma at rest $u^{\mu}=(1,\bm{0})$ (i.e. $\chi=1$), evaluating Eqs. (7) and (8) and assuming $m_{\ell}\approx 0$ yields
$\lambda_{\theta}\;=\;\frac{(1-4\xi)\big{(}\rho_{{}_{\rm T}}-\rho_{{}_{\rm L}}\big{)}}{\rho_{{}_{\rm L}}+(1+4\xi)\rho_{{}_{\rm T}}}\;\simeq\;\frac{\rho_{{}_{\Delta}}}{\rho_{{}_{\rm T}}+\rho_{{}_{\rm L}}}\;,$ (9)
which is displayed in Fig. 1.
Before embedding the fully differential rates on a hydrodynamical background, it is instructive to analyze Eq. (9) for small and large invariant masses, omitting flow conditions. Because the spectral functions are controlled by $M/T$, the large $M$ limit can be viewed as $T\to 0$ for a fixed $M$. The thermal medium becomes irrelevant from the perspective of a highly virtual photon. In this case, $\gamma^{*}$ behaves as it was a massive vector field randomly produced in vacuum, having no preference in its orientation. Hence $\lambda_{\theta}\to 0$ for both LO and NLO in the large $M$ limit, as we shall see below. This asymptotic behaviour is indeed confirmed more quantitatively by the operator product expansion Caron-Huot (2009); Laine et al. (2010), where it was found that $\rho_{{}_{\Delta}}\sim k^{2}(T/M)^{4}$ while $\rho_{{}_{\rm T,L}}\sim M^{2}$ for $M\gg T,k$. For the opposite limit $M/T\to 0$, the virtual photon is more akin to a real photon, lacking the longitudinal polarization, so that $\lambda_{\theta}\simeq(\rho_{{}_{\rm T}}-\rho_{{}_{\rm L}})/(\rho_{{}_{\rm T}}+\rho_{{}_{\rm L}})\to 1$. This is indeed observed at NLO, as revealed by Fig. 1 where, for illustration, we evaluate the spectral functions at $k=k_{\rm av}(M)$ and $\omega=\sqrt{M^{2}+k^{2}}$, with
$k_{\rm av}^{2}(M)\;\equiv\;\frac{\int_{0}^{\infty}{\rm d}k\,k^{4}\exp(-\frac{\sqrt{M^{2}+k^{2}}}{T})}{\int_{0}^{\infty}{\rm d}k\,k^{2}\exp(-\frac{\sqrt{M^{2}+k^{2}}}{T})}\;=\;\frac{3MT\,K_{3}(\frac{M}{T})}{K_{2}(\frac{M}{T})}\;,$ (10)
$K_{n}$ being the modified Bessel function of the second kind. From Fig. 1, we note that the limit for $M\ll T$ is robust even when varying $\alpha_{s}\,$. This physics is absent from LO calculations as the process $q\bar{q}\to\gamma$ is kinematically forbidden, making the LO incapable of converging to the real-photon limit. Therefore the NLO correction is essential for studying polarization, especially in LMR, but also in the IMR. The rest of our study will use $\alpha_{s}=0.3$, a value which is consistent with the bulk of phenomenological studies performed for collisions performed at the LHC Arslandok et al. (2023).
Thermal Dilepton Phenomenology. The relativistic event-by-event heavy-ion collisions of Pb+Pb at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV are simulated using the iEBE-MUSIC framework Schenke et al. (2010, 2011); Paquet et al. (2016), adopting the same model setup as in Ref. Wu et al. (2024). The production of lepton pairs is calculated as it was in Ref. Wu et al. (2024). The local temperature $T(X)$, and local flow velocity $u^{\mu}(X)$ for each fluid cell at space-time point $X=(t,\bm{x})$ are obtained to yield the polarization observables (at LHC energies, the baryon chemical potential $\mu_{B}$ can be neglected). For a given lab-frame virtual photon four-momentum $P$, the final $\lambda_{\theta}$ is a weighted average over all fluid cells, namely
$\lambda_{\theta}(P)\;=\;\frac{\int{\rm d}^{4}X\,\lambda_{\theta}\big{(}P;T(X)\big{)}{\cal R}\big{(}P;T(X)\big{)}}{\int{\rm d}^{4}X\,{\cal R}\big{(}P;T(X)\big{)}},$ (11)
where ${\cal R}\big{(}P;T\big{)}\equiv\frac{{\rm d}R_{\ell\bar{\ell}}}{{\rm d}^{4}K}\big{(}K;T\big{)}\big{/}(1+\lambda_{\theta}\big{(}P,T)/3\big{)}$ is the weight function obtained from integrating over ${\rm d}\Omega_{\ell}$ of the both sides of Eq. (6), and $K^{\mu}=\Lambda^{\mu}_{\nu}(X)P^{\nu}$ is the virtual photon four-momentum in the local rest frame of the fluid cell at $X$, boosted according to the local $u^{\mu}(X)\,$. The invariant mass and transverse momentum, dependent $\lambda_{\theta}(M)$ and $\lambda_{\theta}(p_{{}_{\rm T}})$, are obtained in a similar way with complementary kinematics integrated out.
Figure 2 shows the anisotropy coefficient $\lambda_{\theta}$ as a function of the invariant mass $M$ in the 0-20% centrality class for Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV. Notably, in the LMR the inclusion of NLO corrections changes the anisotropy coefficient qualitatively: $\lambda_{\theta}(M)$ shifts from a near-zero negative to a sizeable positive value after considering the NLO correction (this feature is also evident in Fig. 1). Physically, this indicates that $q\bar{q}$ annihilation tends to produce longitudinal polarization in the HX frame, while the gluon Compton scatterings are more likely to result in $\gamma^{*}$ with transverse polarization.
Figure 2: The anisotropy coefficient $\lambda_{\theta}$ of dilepton polarization with LO and NLO, induced by various sources, as a function of invariant mass $M$ in the 0-20% centrality at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV Pb+Pb collisions at the LHC. The inset figure shows the LO results scaled for clarity. Here, and in the rest of this letter, dilepton results are for dielectrons.
We note that $\lambda_{\theta}$ given in Eq. (7), assuming $m_{\ell}=0$, satisfies $-\frac{1}{3}\leq\lambda_{\theta}|_{M\to 0}=\frac{1-3\chi}{\chi-3}\leq 1\,$. The upper bound is saturated for $\chi=1$ (i.e. $\bm{u}_{*}=0$), as consistent with the behaviour of $\rho_{{}_{\Delta}}/(\rho_{{}_{\rm T}}+\rho_{{}_{\rm L}})$ in Fig. 1. The lower bound occurs when $\chi=0$ (i.e. $\bm{u}_{*}\cdot\hat{k}=0$), giving $\lambda_{\theta}=\frac{-\rho_{{}_{\Delta}}}{\rho_{{}_{\rm L}}+3\rho_{{}_{\rm T}}}\to-\frac{1}{3}$ for $M=0\,$. Therefore, once local flow conditions are included, $\lambda_{\theta}$ at $M\simeq 0$ limit will be slightly less than 1, as realized in Fig. 2 for the NLO result. Nevertheless, although diluted by the plasma kinematics, the results obtained with NLO and LO rates are still radically different in the limit $M\to 0$, since the LO spectral function fails to capture the relevant physics in the IMR. In the HMR, QGP dileptons become almost unpolarized as $\lambda_{\theta}\approx 0$, consistent again with the pure thermal prediction in Fig. 2. However, the Drell-Yan process, the consideration of which lies beyond the scope of this study, is the main source of dilepton production and polarization in the HMR. In addition, Fig. 2 shows the effect on polarization of a pre-equilibrium contribution calculated with KøMPøST as described in Ref. Gale et al. (2022). That pre-hydro addition, included in the “total” polarization, has a noticeable but relatively modest effect here.
In Fig. 3, we show $\lambda_{\theta}$ as a function of transverse momentum $p_{{}_{\rm T}}$ in the LMR and the IMR, where contributions from different stages of the fireball evolution are also displayed. In a given range of invariant mass, $\lambda_{\theta}$ increases as a function of $p_{{}_{\rm T}}$. The behavior is consistent with that observed earlier: as the transverse momentum grows, the influence of the invariant mass shrinks, bringing the kinematics closer to those of real photons. In that limit, the transverse polarization dominates over the longitudinal. In the fluid dynamical environment, the exact value is also affected by the averaging over the flow. For the opposite limit $p_{{}_{\rm T}}\to 0$, we observe that $\lambda_{\theta}$ does not vanish, which can also be attributed to the non-zero flow velocity of the medium in the HX frame of the dilepton.
Figure 3: The anisotropy coefficient $\lambda_{\theta}$ of dilepton polarization evaluated at NLO, induced by various sources, as a function of transverse momentum $p_{{}_{\rm T}}$ in the LMR (panel (a)) and IMR (panel (b)), for the 0-20% centrality class in Pb + Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV.
Since gluon-mediated processes only contribute to dilepton polarization at NLO (and beyond), the corresponding enhancement of $\lambda_{\theta}$ in the IMR window, if experimentally confirmed, could be taken as a flag for such elementary QCD interactions. Moreover, the behaviour of $\lambda_{\theta}(p_{{}_{\rm T}})$ could indicate the relative abundance of quarks and gluons: a larger $\lambda_{\theta}(p_{{}_{\rm T}})$ means a greater gluon-quark ratio during the pre-equilibrium stage than that in thermal equilibrium, as is clear in the bottom panel of Fig. 3. It is therefore very promising that the $p_{{}_{\rm T}}$ dependence of $\lambda_{\theta}$ in the IMR will offer valuable insights into the pre-equilibrium stage of heavy-ion collisions in future studies.
Summary. In this letter, we perform the first study of dilepton polarization phenomenology using multi-stage heavy-ion collision simulations for $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV Pb+Pb collisions at the LHC. We employed NLO thermal emission rates in our analysis, which were coupled with the local temperature and flow velocity provided by the complete evolution history of the heavy-ion simulation.
We observe an overall sign change and a significant enhancement in $\lambda_{\theta}$ when going from LO to NLO. This better matches the physical expectation that at small invariant mass, the virtual photon behaves like a real photon with two transverse polarization modes. In contrast, the LO calculation does not capture this aspect due to the vanishing phase space for $M\to 0\,$, where the NLO result is actually the “leading” contribution. At large invariant masses, the NLO result is a small correction to the LO one and both reflect the lack of a preferred polarization direction for $M\gg T$. Underlying this dramatic difference between orders in the perturbative calculation, is the participation of (thermal) gluons in the production reactions. The importance of gluons is readily supported from non-perturbative lattice studies, where the current-current correlators do not change qualitatively when going from quenched to $2$+$1$ flavour QCD Brandt et al. (2018); Cè et al. (2022); Ali et al. (2024).
We also explore the dilepton polarization phenomenon induced during the pre-equilibrium stage and the QGP phase. It is found that dilepton polarization can be highly sensitive to the pre-equilibrium evolution; we find that the effects of the pre-equilibrium phase on dilepton polarization become more pronounced when considering the transverse momentum $p_{{}_{\rm T}}$-dependence in the IMR. Therefore, studies of the pre-equilibrium stages in heavy-ion collisions will benefit from those of dilepton polarization.
Acknowledgements. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) [SAPIN-2020-00048 and SAPIN-2024-00026], and in part by US National Science Foundation (NSF) under grant number OAC-2004571. G.J. was funded by the Agence Nationale de la Recherche (France), under grant ANR-22-CE31-0018 (AUTOTHERM). Computations were made on the Béluga super-computer system from McGill University, managed by Calcul Québec and by the Digital Research Alliance of Canada.